# Learning-guided Kansa collocation for forward and inverse PDEs beyond linearity

### v1

@@ 1 Introduction
It’s unknown how (extended) CNF solvers Zhong et al. (2023) compared with other classical and neural PDE solvers. Hence, we also implement and evaluate selected prior methods on the benchmarks on their effectiveness with different quality metrics (e.g. L1, L2, errors) against ground truth solutions, efficiency, computation resource, convergence speed, method complexity, and finally utility in research, i.e. their scientific simulation applications or integration with other methods, e.g. differentiable rendering to solve inverse physics-related problems in Graphics.
@@ 3.2 Kansa collocation
Kernel functions. Radial Basis Function (RBF) inversely relates the distance $r$ between the input $\mathbf{x}$ and a fixed origin point $\mathbf{c}$ to the output value.
@@ 3.2.1 Linear operator case
$\mathcal{F}[\hat{u}](\mathbf{x}_{i})=\mathcal{F}[\sum_{i=1}^{N}\alpha_{i}\cdot\psi_{i}](\mathbf{x}_{i})=\sum_{i=1}^{N}\alpha_{i}\cdot\mathcal{F}[\psi_{i}](\mathbf{x}_{i})=h(\mathbf{x}_{i}).$ (7)
@@
$\psi_{i}=e^{-\frac{r_{i}^{2}}{2\sigma^{2}}},\quad r_{i}^{2}=(x-x_{i})^{2}+(t-t_{i})^{2}.$ (9)
@@
$\mathcal{F}[\psi_{i}](\mathbf{x}_{j})=\frac{\partial\psi_{i}(\mathbf{x}_{j})}{\partial t}=\frac{\partial\psi_{i}(\mathbf{x}_{j})}{\partial r_{i}^{2}}\cdot\frac{\partial r_{i}^{2}}{\partial t}=-\frac{1}{2\sigma^{2}}e^{-\frac{r_{i}^{2}}{2\sigma^{2}}}\cdot 2(t_{j}-t_{i})=-\frac{t_{j}-t_{i}}{\sigma^{2}}e^{-\frac{r_{i}^{2}}{2\sigma^{2}}}.$ (10)
@@ 3.2.3 Extension 2: nonlinear operator case
We can invert equation 5 $\mathbf{a}=\mathbf{K}^{-1}\cdot\mathbf{u}$, assuming $\mathbf{K}$ invertibility from independent basis. By substituting $\mathbf{a}$ into $\mathbf{u}^{\prime}=\mathbf{K_{x}}\cdot\mathbf{a}$, one gets $\mathbf{u}^{\prime}=\mathbf{K_{x}}\cdot\mathbf{K}^{-1}\cdot\mathbf{u}$. The differentiable matrix is thus,
@@ 3.2.4 Auto-tuning of Kansa hyperparameters
To tune the key Kansa method hyperparameter, kernel shape parameter $\epsilon$ in equation 3, Zhong et al. (2023) proposed one of the self-tuning methods for $\epsilon$, by minimizing the variation of the solution field $u$ over all collocation points, and the condition number of the kernel matrix $\mathbf{F}$,
@@
For non-linear operator case, the solution $u$ depends on $\epsilon$ implicitly via the coefficients $\alpha_{i}$. The matrix $\mathbf{F}$ no longer exists explicitly. Here, we propose to directly minimize the PDE residuals over all collocation points, the total variation of the solution field $u$, and the training L2 loss between the predicted solution $u$ and the ground truth solution $u^{gt}$ if available,
@@
where $\omega_{1},\omega_{2}$, and $\omega_{3}$ are the weights for each penalty term.
@@ 4.1 Performance metrics
Accuracy. Given the numerical solution $\hat{u}_{i}$ from PDE solvers, and the corresponding ground truth $u_{i}$, the $L_{2}$ risk $\mathcal{R}_{L_{2}}$ is the average discretized error over all $N_{\text{test}}$ test points,
@@ 4.3 Extension 1: Kansa method for coupled PDEs
The Lotka-Volterra equations equation 47 are initialized as per Table LABEL:tab:lotka-volterra-setup, where the number of domain quadrature points $N_{\mathcal{R}}=100$, and initial condition points $N_{d}=1$. For evaluation, the test points $N_{\text{test}}=64$. The results from Kansa method are summarized in Table LABEL:tab:LV-maxwell-error, where the Gaussian RBF shape parameters, as defined in equation 3, are set as $\epsilon=0.2$ for both $x(t)$ and $y(t)$.
@@
Table 6: 1D Maxwell’s equations experimental setup.
domain time span parameter initial conditions
@@ 4.3.1 Forward problem
Accuracy. The results from Kansa method are summarized in Table LABEL:tab:LV-maxwell-error. Both errors converge with increasing $C_{\text{scale}}$. Efficiency. The training time and inference time of Kansa method on Lotka-Volterra equations are $0.4034$ and $0.0001$ seconds, respectively. The training time and inference time of Kansa method on Maxwell’s equations are $0.4486$ and $0.0005$ seconds.
@@ 4.4.1 Forward problem
$C^{t}_{\text{scale}}$ 1 2 4 10
[…]
Stability unstable unstable stable stable
@@ Appendix D Environment setup
Appendix E Learning theory
E.1 Functional analysis
We introduce basic functional analysis concepts here for PDE solvers and later learning theory (Appendix § E).
@@ E.1 Functional analysis
For additional concepts used for learning theory, please refer to Appendix § E.2.
E.2 Functional analysis addendum
@@
E.3 Approximation theory of neural networks
@@ E.4 Error analysis
E.5 PINN learning theory

### v2

# arXiv:2602.07970v2[cs.CE] 02 Mar 2026

@@ 1 Introduction
It’s unknown how (extended) Constrained Neural Fields (CNF) Zhong et al. (2023) compared with other classical and neural PDE solvers. Hence, we also implement and evaluate selected prior methods on the benchmarks on their effectiveness with different quality metrics (e.g. L1, L2, errors) against ground truth solutions, efficiency, computation resource, convergence speed, method complexity, and finally utility in research, i.e. their scientific simulation applications or integration with other methods, e.g. differentiable rendering to solve inverse physics-related problems in Graphics.
@@ 3.2 Kansa collocation
Kernel functions. Radial Basis Function (RBF) relates the distance $r$ between the input $\mathbf{x}$ and a fixed origin point $\mathbf{c}$ to the output value.
@@ 3.2.1 Linear operator case
$\mathcal{F}[\hat{u}](\mathbf{x}_{i})=\mathcal{F}[\sum_{k=1}^{N}\alpha_{k}\cdot\psi_{k}](\mathbf{x}_{i})=\sum_{k=1}^{N}\alpha_{k}\cdot\mathcal{F}[\psi_{k}](\mathbf{x}_{i})=h(\mathbf{x}_{i}).$ (7)
@@
$\psi_{k}=e^{-\frac{r_{k}^{2}}{2\sigma^{2}}},\quad r_{k}^{2}=(x_{k}-x_{i})^{2}+(t_{k}-t_{i})^{2}.$ (9)
@@
$\mathcal{F}[\psi_{k}](\mathbf{x}_{i})=\frac{\partial\psi_{k}(\mathbf{x}_{i})}{\partial t}=\frac{\partial\psi_{k}(\mathbf{x}_{i})}{\partial r_{k}^{2}}\cdot\frac{\partial r_{k}^{2}}{\partial t}=-\frac{1}{2\sigma^{2}}e^{-\frac{r_{k}^{2}}{2\sigma^{2}}}\cdot 2(t_{k}-t_{i})=-\frac{t_{k}-t_{i}}{\sigma^{2}}e^{-\frac{r_{k}^{2}}{2\sigma^{2}}}.$ (10)
@@ 3.2.3 Extension 2: nonlinear operator case
By inverting equation 5, $\mathbf{a}=\mathbf{K}^{-1}\cdot\mathbf{u}$, assuming $\mathbf{K}$ invertibility from independent basis. By substituting $\mathbf{a}$ into $\mathbf{u}^{\prime}=\mathbf{K_{x}}\cdot\mathbf{a}$, one gets $\mathbf{u}^{\prime}=\mathbf{K_{x}}\cdot\mathbf{K}^{-1}\cdot\mathbf{u}$. The differentiable matrix is thus,
@@ 3.2.4 Auto-tuning of Kansa hyperparameters
To tune the key Kansa method hyperparameter, kernel shape parameter $\epsilon$ in equation 3, Zhong et al. (2023) proposed one of the self-tuning methods for $\epsilon$ by minimizing the variation of the solution field $u$ over all collocation points, and the condition number of operator-evaluated kernel matrix $\mathbf{F}$,
@@
For non-linear operator case, the solution $u$ depends on $\epsilon$ implicitly via the coefficients $\alpha_{i}$. The matrix $\mathbf{F}$ no longer exists explicitly. Here, we propose to directly minimize the PDE residuals over all collocation points, the total variation of the solution field $u$, and the training L2 loss between the predicted solution $u$ and the ground truth solution $u^{gt}$ if training data are available,
@@
where $\omega_{1},\omega_{2}$, and $\omega_{3}$ are the penalty weights. Grid search is used as an optimizer.
@@ 4.1 Performance metrics
Accuracy. Given the numerical solution $\hat{u}_{i}$ from PDE solvers, and the ground truth $u_{i}$, the $L_{2}$ risk $\mathcal{R}_{L_{2}}$ is the average discretized error over all $N_{\text{test}}$ test points on the spatial-temporal domain,
@@ 4.3 Extension 1: Kansa method for coupled PDEs
The Lotka-Volterra equations equation 47 are initialized as per Table LABEL:tab:lotka-volterra-setup, where the number of domain quadrature points $N_{\mathcal{R}}=100\times 1$, and initial condition points $N_{d}=1$. For evaluation, the test points $N_{\text{test}}=64$. The results from Kansa method are summarized in Table LABEL:tab:LV-maxwell-error, where the Gaussian RBF shape parameters, as defined in equation 3, are set as $\epsilon=0.2$ for both $x(t)$ and $y(t)$.
@@ 4.3.1 Forward problem
Accuracy. The results from Kansa method are summarized in Table LABEL:tab:LV-maxwell-error. Both errors converge with increasing $C_{\text{scale}}$ as defined above. Efficiency. The training time and inference time of Kansa method on Lotka-Volterra equations are $0.4034$ and $0.0001$ seconds, respectively. The training time and inference time of Kansa method on Maxwell’s equations are $0.4486$ and $0.0005$ seconds.
@@ Appendix D Environment setup
Table 14: 1D Maxwell’s equations experimental setup.
domain time span parameter initial conditions
Appendix E Burgers’ equation stability experiment
$C^{t}_{\text{scale}}$ 1 2 4 10
[…]
Stability unstable unstable stable stable
Appendix F Learning theory
F.1 Functional analysis
We introduce basic functional analysis concepts here for PDE solvers and later learning theory (Appendix § F).
@@ F.1 Functional analysis
For additional concepts used for learning theory, please refer to Appendix § F.2.
F.2 Functional analysis addendum
@@
F.3 Approximation theory of neural networks
@@ F.4 Error analysis
F.5 PINN learning theory